Solve this equation using the square root property. If possible, simplify radicals.All answers should be exact.(n−7)2=51
The solution(s) to this equation have the form n=A±√B

Found 2 solutions by MathLover1, mananth:
Answer by MathLover1(20849)
Answer by mananth(16946)

(n−7)^2 = 51
(n-2) = (+/-)(sqrt(51))
n = 2 +/- sqrt(51)